# Agent A/B: Automated and Scalable Web A/B Testing with Interactive LLM Agents

### v3

@@ 4. Agent A/B
Based on feedback from the formative study, we designed and developed Agent A/B a modular system that enables automated A/B testing on live websites with LLM agents to simulate realistic and diverse user interactions. Traditional A/B testing is bottle-necked by a number of long-standing limitations, including the expensive requisites of large sample sizes and user traffic for significant results, the complicated and prolonged procedures for experiment preparation, execution, and analysis, as well as the limited resources that can only support experiments with very few designs/features. However, our system takes advantage of LLM agents for low-cost, rapid feedback loop, and high-throughput evaluation of interface variants without relying on live user deployment. The system is designed with modularity and extensibility in mind, which can be accommodated in different web environments, target user populations, and LLM agent architectures.
In this section, we describe the Agent A/B system architecture in detail, including the end-to-end A/B testing pipeline, the interaction loop of LLM agent and the web environment, and the implementation techniques that support robust and high-fidelity simulation possible on live web platforms.
@@ 4.1. System Overview and Pipeline
Agent A/B is designed as an end-to-end simulation system for LLM agent-based A/B testing in live browser environments. Analogous to the preparation stage of real user A/B testing, the end users of our system (i.e., UX researchers or product managers) need to determine details of A/B testing designs and provide two web environment variants to be tested. After taking the user inputs, the system operates on four LLM-powered modules (Fig. 1): LLM agent generation, testing preparation, autonomous A/B testing simulation, and post-testing analysis.
Once these inputs are provided by end users, the LLM Agent Generation Module queries the backend LLM to generate the specified number of LLM agents with diverse persona descriptions and intentions. The query explicitly instructs the LLM to ensure the generated LLM agent persona and intentions must follow the user-provided agent specifications. In our experiment in Section 5, we leveraged this module to generate 100,000 agents. We provide a sample persona generated by our system in Appendix A. After generating the LLM agents as the candidate A/B testing participants, the Testing Preparation Module performs the agent traffic allocation by splitting the agents into control (without new feature) and treatment (with new feature) groups, and each group is assigned to interact with the corresponding web environments. The statistics of agent characteristic distributions will be calculated within each group to ensure that the distribution of the LLM agents is relatively balanced; if the statistics are not balanced, the Testing Preparation Module will re-execute agent traffic allocation until satisfactory.
These web environments for both groups need to be launched using independent browser instances controlled through ChromeDriver (for web environment parsing) and Selenium WebDriver (for automated interaction execution) integration. During the automatic interaction with the web environment, each LLM agent begins to interact with the webpage using an autonomous action prediction loop, which is shown in Fig. 3. Each loop involves perceiving the current webpage state, interpreting the action space, predicting the next action, and executing that action in the browser. The details of this core automated agent-environment interaction process are explained below in Section 4.2. During the process, each step of the interaction is recorded, and the system monitors the overall session progression until termination, such that the agent accomplishes the intended goal or encounters failure cases.
Figure 3. One action prediction iteration of the automated web testing in Agent A/B with an LLM agent. (1) An Agent Profiling Module maintains a comprehensive agent description with an LLM-generated persona, user-specified intention, and the action history of the current session. In the meantime, (2) the Environment Parsing Module parses the webpage into structured web representation and action spaces. (3) All the information is fed into the LLM Agent for the next action prediction, (4) which will be executed by the Action Execution Module in the web environment to drive the next iteration. :
4.2. Agent-Environment Interaction Architecture
At the core of Agent A/B is an iterative mechanism where each LLM agent continuously interacts with a real web environment by dynamically updating its understanding of the environment and adjusting its actions accordingly. This architecture consists of three tightly integrated components: the Environment Parsing Module, the LLM agent, and the Action Execution Module. Together, these components enable robust operation in complex and dynamic web environments, for instance, our system evaluation in Section 5 is experimented on the Amazon [^1] http://www.Amazon.com platform.
Environment Parsing Module.
LLM Agent.
The LLM agent functions as a decision-making module that consumes the current state and outputs the next action to be taken. In particular, the LLM agent models the next-step decision-making problem as a form of language-based reasoning and planning task by mapping structured state observations into reasoning traces and action predictions. Agent A/B is not bound to a specific LLM agent implementation. Instead, our system treats the LLM agent as an exchangeable module that supports various types of LLM web agents (ReAct (Yao et al., 2023), FireClaw [^2] https://github.com/mendableai/firecrawl-mcp-server) with convenient “plug-and-play” APIs, analogous to the Model Context Protocol (MCP) proposed by Claude. In our experiment, we adopt the UXAgent framework (Lu et al., 2025b), a state-of-the-art LLM agent for web interactions, as a representative implementation due to its strong performance and support for multi-step planning and intermediate memory.
Figure 4. The two design variants of the left filter panel on Amazon.com for the A/B testing case study with Agent A/B. In the control condition (left screenshot), all the filter options are shown; in the treatment condition (right screenshot), we apply a similarity-based ranking algorithm to reduce the filter options that have lower than 80% similarity to the users’ search query (e.g., “solar filter for telescope”).:
Action Execution Module.
5. Case Study: A/B Testing with Reduced Filter Options on Amazon.com
To evaluate the effectiveness of Agent A/B, we conducted large-scale simulations on a live e-commerce platform.
5.1. Method: A/B Testing Scenario
We designed a simulated A/B testing scenario focused on a novel feature design of the left-side filtering panel of Amazon’s shopping interface, a common area for UI experimentation in e-commerce. The goal was to evaluate how variations in filter configurations influence user behaviors such as dollar spending and product purchases.
@@ 5.1. Method: A/B Testing Scenario
Table 2. Comparison of human customers’s actions per session in control condition (current design with full filter list), and the LLM agents as virtual customers in both control condition and treatment condition (new design with reduced filter list). The LLM agents in treatment condition purchased significantly more items than the ones in the control condition. $\chi^{2}(1)=5.51$, p-value ¡ 0.05
@@
Our team has the privilege of altering the left filter panel interface (Fig. 4). In comparison to the existing design as the control condition, where all the filter options are shown, the new design (treatment condition) uses a similarity-based ranking algorithm to reduce the filter options that have a lower-than-80% similarity score to the users’ search query.
@@
Using the LLM Agent Generation Module, we created 100,000 agent personas [^3] We will open source these agent personas and randomly sampled 1,000 to simulate individual shopping sessions, 500 assigned into the treatment condition and the other 500 assigned into the control condition. In parallel, we also conducted the real human online A/B testing with 1 million users randomly assigned into the control condition and another 1 million users in the treatment condition. Each agent was initialized with a persona profile and a shopping task (e.g., “find a budget smart speaker under \$40 with strong customer reviews”). The allowed action space for the LLM agent includes typical e-commerce flows: search, filter use, product examination, and cart actions. Persona generation followed the methodology in Chen et al. (2024) with LLM agent’s demographics (age, income, occupation), preferences, and shopping goals.
@@
5.2. Finding: Alignment with Human Behavior
To benchmark our simulation system in aligning with human behavior, we conducted a simulated A/B testing case study using Agent A/B and also have access to the results of online A/B testing results with real human subjects using identical task conditions. Table 2 presents the aggregated actions and statistics from sessions with the 1,000 LLM agents participating and with 2M human participants [^4] Due to company policy, the human participant’s behavior data in the treatment condition can not be shared..
As shown in Table 2, human participants in the control condition engaged in longer, more exploratory interactions than agents, averaging nearly twice as many actions per session (15.96 vs. 6.05). They also clicked on more product detail pages (6.96 vs. 1.87) and performed more searches (6.40 vs. 1.42). Despite these differences in interaction style, humans and agents showed similar purchase rates (0.62 vs. 0.81) and comparable use of filter options (0.33 vs. 0.58).
These findings highlight notable behavioral differences: while LLM agents adopt more goal-directed, structured interaction strategies, human users display broader exploratory behaviors. Nevertheless, the decision-making processes and general interaction patterns of the agents reasonably approximate human behavior for intention-driven tasks. This alignment supports the utility and validity of employing agent simulations for controlled UX evaluations, especially in scenarios where rapid interface assessment without live deployment is beneficial.
Figure 5. Number of LLM agents who completed a purchase under control and treatment conditions.:

### v4

@@ 4.1. System Overview and Pipeline
Agent A/B is an end-to-end system for simulating A/B tests on websites with persona-driven LLM agents. Experiment owners provide (1) agent specifications (e.g., number of agents and persona attributes), (2) A/B testing configurations (e.g., task intentions and behavioral metrics), and (3) two functional web variants for the control and treatment conditions (Fig. 1). Given these inputs, Agent A/B orchestrates four coordinated modules: (i) LLM Agent Generation, (ii) Testing Preparation, (iii) Autonomous A/B Simulation, and (iv) Post-Testing Analysis.
First, the LLM Agent Generation Module instantiates a population of agents by generating diverse personas and pairing them with the user-specified task intentions, ensuring variability in demographics and behavioral tendencies while adhering to the experiment constraints. The experiment owner specifies a target demographic distribution and provides an example persona as a style reference; a persona pool is initialized with that example. The system then iteratively samples an existing persona from the pool, samples demographic attributes from the target distribution, and prompts the LLM to produce a new persona matching both the sampled demographics and the reference style. Each generated persona is added back to the persona pool, enabling diverse yet stylistically consistent personas at scale. Second, the Testing Preparation Module assigns agents to control and treatment conditions via traffic allocation and checks balance across key persona attributes, re-splitting when needed to reduce distributional skew. Third, in the Autonomous A/B Simulation Module, agents interact with their assigned web variant in isolated browser sessions using a recurrent perceive–decide–act loop (Section 4.2), while the system records step-level traces of agent actions and web states. Finally, the Post-Testing Analysis Module aggregates traces across conditions to compute outcome metrics (e.g., completion or purchase rates), interaction efficiency measures (e.g., actions per session, session duration), and behavior patterns (e.g., search or filter usage), and supports stratified analyses by persona attributes to surface subgroup differences. Together, these modules enable scalable comparative evaluation of interface variants and provide early behavioral signals without relying on live user traffic. Full details of the system design can be found in Appendix A.
@@
Table 1. Comparison of human customers’ actions per session in control condition (current design with full filter list), and the LLM agents as virtual customers in both control condition and treatment condition (new design with reduced filter list). The LLM agents in treatment condition purchased significantly more items than the ones in the control condition. $\chi^{2}(1)=5.51$, p-value ¡ 0.05
4.2. Agent-Environment Interaction Architecture
At the core of Agent A/B is an iterative interaction loop in which each LLM agent operates directly on a live web environment and continuously adapts its actions based on evolving page state. This loop consists of three tightly coupled components: an Environment Parsing Module, an LLM Agent, and an Action Execution Module (Fig. 3). Together, they enable robust and scalable interaction with dynamic, real-world webpages.
At each step, the Environment Parsing Module extracts the current webpage into a structured representation of task-relevant elements (e.g., products, filters, prices) and the available action space, reducing noise from raw HTML or visual clutter. This observation, together with the agent’s persona, intention, and interaction history, is passed to the LLM Agent, which performs language-based reasoning and planning to select the next structured action (e.g., search, click filter, select item, purchase), independent of any specific agent architecture. The Action Execution Module then translates this action into browser-level interactions and executes it in the live environment, incorporating basic fault handling such as retries or re-parsing to accommodate dynamic page changes. This perceive–decide–act loop continues until task completion or failure, while Agent A/B logs fine-grained traces of page states, actions, and outcomes for downstream A/B-style analysis.
5. Case Study: A/B Testing with Reduced Filter Options on Amazon.com
We evaluate Agent A/B through a simulated A/B testing case study on Amazon’s shopping interface, focusing on a redesign of the left-side filtering panel, a common target for UI experimentation in e-commerce. The treatment condition replaces the full filter list with a similarity-based ranking design that suppresses low-relevance options, while the control condition retains the existing interface. Using Agent A/B, we generated a large pool of agent personas and sampled 1,000 agents (500 per condition) to simulate shopping sessions under realistic demographic distributions and task intentions, in parallel with a large-scale human A/B test on the same feature. Agents interacted with the live website using standard e-commerce actions, and the system logged full interaction traces and outcomes for analysis. Full details of the scenario design, agent setup, infrastructure, and cost analysis are provided in Appendix B.
5.1. Finding: Alignment with Human Behavior
To assess how well Agent A/B aligns with real user behavior, we compared a simulated A/B test using 1,000 LLM agents with results from a parallel online A/B experiment involving 2M human participants under identical task conditions (Table 1). Humans exhibited more exploratory interaction patterns, including longer sessions and more frequent searches and product views, while agents followed more goal-directed trajectories with fewer actions. Despite these differences in interaction style, agents and humans showed comparable purchase rates and similar use of filter options, indicating alignment in decision outcomes for intention-driven tasks. Overall, these results suggest that while LLM agents simplify exploratory behavior, they capture key decision-making signals relevant for early-stage UX evaluation, supporting the use of agent-based simulation as a complementary tool to human A/B testing.
5.2. Findings: System Effectiveness Across Interface Variants
To evaluate whether Agent A/B can detect differences between interface variants, we compared LLM agent behavior across control and treatment conditions with different filter designs. As shown in Fig. 6 and Table 1, agents in the treatment condition interacted more with products and filters and completed more purchases than those in the control condition (414 vs. 404; $\chi^{2}(1)=5.51$, $p=0.03$), indicating a modest but statistically reliable increase. Average spending exhibited a small upward trend under the treatment design, though this effect did not reach statistical significance.
Subgroup analyses further revealed heterogeneous responses across personas (Fig. 6): agents representing older and male customers showed larger increases in spending under the simplified filter design, while younger agents exhibited decreased spending, suggesting differing preferences for filter complexity. Importantly, these agent-based patterns aligned directionally with results from human A/B testing, supporting Agent A/B ’s ability to surface meaningful behavioral signals and detect subtle interface effects prior to live deployment.
6. Discussion
Our design and evaluation of Agent A/B demonstrate how agent-based simulation can serve as a design support mechanism in the web interface design lifecycle. Rather than replacing real-user experimentation, Agent A/B complements existing workflows by enabling earlier, faster, and lower-risk exploration, aligning with broader trends in HCI that leverage large-scale simulation for iterative design and decision-making (Park et al., 2023; Horton, 2023; Gao et al., 2024).
Simulation-Supported Design Iteration and Exploration.
A key contribution of Agent A/B is its ability to accelerate design iteration by enabling agent-based pilot experiments prior to A/B testing. Although A/B testing remains the gold standard for design validation, it is constrained by limited user traffic, long feedback cycles, and substantial engineering and organizational overhead (Fabijan et al., 2017; Kohavi et al., 2013; Xu et al., 2015). By supporting rapid evaluation of interface variants through agent simulations, Agent A/B allows designers to obtain early signals without implementing and deploying their new design to real users. This capability extends prior prototyping and experimentation tools (Lasecki et al., 2015; Hartmann et al., 2006; Kuznetsov et al., 2022) by grounding early feedback in end-to-end interaction traces, and enables simulation-driven methodologies such as parallel evaluation and computational design exploration (Kumar et al., 2013).
Inclusive Piloting and Risk-Aware Testing.
Beyond accelerating iteration, Agent A/B supports inclusive and risk-aware piloting for user populations that are difficult, costly, or ethically sensitive to recruit. Groups such as older adults or users with lower digital literacy are often underrepresented in early experimentation, even though design flaws may disproportionately affect them. By configuring agent personas to reflect diverse demographic and behavioral characteristics, designers can explore how features may impact different user groups before exposing real users to potential harm. This capability complements prior crowd-based and proxy evaluation approaches (Bernstein et al., 2011; Oppenlaender et al., 2020) by offering an automated alternative that helps surface subgroup disparities early, enabling more robust and informed A/B testing, and aligns with emerging evaluation techniques using agent judges (Chen et al., 2025c).
7. Conclusion
In this paper, we presented Agent A/B, a system that enables large-scale, LLM agent-based simulation of A/B testing for web interfaces. Our evaluation demonstrated that LLM agents exhibit realistic, goal-aligned behaviors, and are sensitive to interface variations. By supporting rapid, risk-free behavioral piloting, Agent A/B introduces a new phase for agent-based piloting in the design life-cycle that complements traditional A/B testing and expands the scope of early-stage UX evaluation. We envision future extensions that further enhance agent fidelity, broaden domain coverage, and integrate simulation into intelligent design optimization workflows.
Appendix A Agent A/B
Figure 2. The workflow of web A/B testing and three challenges reported from formative study: (1) the cost and difficulty of securing considerable user traffic for significant results, (2) the whole A/B testing period can span across weeks and months, and (3) limited testing opportunites.
A timeline diagram illustrates the standard A/B testing workflow and highlights three key challenges. The timeline includes seven sequential phases: Feature Ideation (new UI layout or functionalities), A/B Testing Experiment Design (testing configuration and implementation), Data Monitoring (collect user interactions and rationales), Feature Deployment (final design decisions), Post-Testing Analysis (behavior patterns, failure cases, outcomes), Traffic Allocation (real user traffic toward variants), and Team Alignment (decision maker approval or rejection). The A/B Testing Period is marked across the Data Monitoring and Traffic Allocation stages. Three challenges are highlighted in red: (1) Scalable recruitment difficulties, (2) Limited testing opportunities, and (3) Slow user feedback loop. Each stage is connected along a central purple timeline.
Figure 3. One action prediction iteration of the automated web testing in Agent A/B with an LLM agent. (1) An Agent Profiling Module maintains a comprehensive agent description with an LLM-generated persona, user-specified intention, and the action history of the current session. In the meantime, (2) the Environment Parsing Module parses the webpage into structured web representation and action spaces. (3) All the information is fed into the LLM Agent for the next action prediction, (4) which will be executed by the Action Execution Module in the web environment to drive the next iteration. The light-weight Environment Parsing Module and Action Execution Module are web-site specific by their nature, and other modules include LLM Agent are generalizable to different websites.
A system diagram shows how a language model-based agent interacts with a web environment. The process begins with a web interface (Control/Treatment) feeding into an Environment Parsing Module, which extracts the webpage as a JSON structure and defines an action space (e.g., search, filter, purchase, product_click, terminate). This information, along with a predefined persona and intention, is sent to the LLM Agent (Action Prediction), which outputs the next action (e.g., search “solar filter for telescope”). This action is executed through the Action Execution Module, which sends it back to the Web Environment, completing the interaction loop.
A.1. System Overview and Pipeline
Agent A/B is designed as an end-to-end simulation system for LLM agent-based A/B testing in live browser environments. Analogous to the preparation stage of real user A/B testing, the experiment owners (i.e., UX researchers or product managers) need to determine details of A/B testing designs and provide two web environment variants to be tested. After taking the user inputs, the system operates on four LLM-powered modules (Fig. 1): LLM agent generation, testing preparation, autonomous A/B testing simulation, and post-testing analysis.
Once these inputs are provided by the experiment owners, the LLM Agent Generation Module queries the backend LLM to generate the specified number of LLM agents with diverse persona descriptions and intentions. The query explicitly instructs the LLM to ensure the generated LLM agent persona and intentions must follow the user-provided agent specifications. In our experiment in Section 5, we leveraged this module to generate 100,000 agents. We provide a sample persona generated by our system in Appendix C. After generating the LLM agents as the A/B testing participants, the Testing Preparation Module performs the agent traffic allocation by splitting the agents into control (without new feature) and treatment (with new feature) groups, and each group is assigned to interact with the corresponding web environments. The statistics of agent characteristic distributions will be calculated within each group to ensure that the distribution of the LLM agents is balanced; if the statistics are not balanced, the Testing Preparation Module will re-execute agent traffic allocation until satisfactory.
These web environments for both groups need to be launched using independent browser instances controlled through ChromeDriver (for web environment parsing) and Selenium WebDriver (for automated interaction) integration. During the automatic interaction with the web environment, each LLM agent begins to interact with the webpage using an autonomous action prediction loop, which is shown in Fig. 3. Each loop involves perceiving the webpage state, interpreting the action space, predicting the next action, and executing that action in the browser. . During the process, each step of the interaction is recorded, and the system monitors the overall session progression until termination, such that the agent accomplishes the intended goal or encounters failure cases.
A.2. Agent-Environment Interaction Architecture
At the core of Agent A/B is an iterative mechanism where each LLM agent continuously interacts with a real web environment by dynamically updating its understanding of the environment and adjusting its actions accordingly. This architecture consists of three tightly integrated components: the Environment Parsing Module, the LLM agent, and the Action Execution Module. Together, these components enable robust operation in complex and dynamic web environments, for instance, our system evaluation in Section 5 is experimented on the Amazon [^1] http://www.Amazon.com platform. We will release our code upon acceptance of this paper.
Environment Parsing Module.
LLM Agent.
The LLM agent functions as a decision-making module that consumes the current state and outputs the next action to be taken. In particular, the LLM agent models the next-step decision-making problem as a form of language-based reasoning and planning task by mapping structured state observations into reasoning traces and action predictions. Agent A/B is not bound to a specific LLM agent. Instead, our system treats the LLM agent as an exchangeable module that supports various types of LLM web agents (ReAct (Yao et al., 2023), FireClaw [^2] https://github.com/mendableai/firecrawl-mcp-server) with convenient “plug-and-play” APIs, analogous to the Model Context Protocol (MCP) proposed by Claude.
In our experiment, we adopt the UXAgent framework (Yuxuan Lu et al., 2025), a state-of-the-art LLM agent for web interactions, as a representative implementation due to its strong performance and its built-in support for multi-step planning, hierarchical control, and intermediate memory management. UXAgent (Fig. 4) follows a dual-loop architecture composed of a Fast Loop for rapid perception–action cycles and a Slow Loop for deeper reasoning. Within the fast loop, the perception module converts raw signals from the browser connector into natural-language observations, which are then appended to the memory stream. The planning module generates short-horizon plans conditioned on these observations, and the action module executes the action. The slow loop complements this by incorporating a wonder module that generates speculative or exploratory thoughts, simulating human-like mind wandering, and a reflection module that performs deliberate reasoning over past trajectories and current goals. Both loops continuously read from and write to the unified memory stream, allowing perception, planning, action, wonder, and reflection to jointly shape the agent’s evolving internal state and decision process.
Figure 4. UXAgent Architecture design (Yuxuan Lu et al., 2025).
The architecture features two interacting processes: the Fast Loop and the Slow Loop, which handle rapid web interactions and deeper reasoning respectively. Within the fast loop, the perception module converts raw observations from the environment into natural language memory entries stored in the memory stream, while the planning module generates short-term plans that the action module executes. In the slow loop, the wonder module simulates human “mind wandering” and the reflection module performs more thorough reasoning about the current context. Modules retrieve relevant entries from the memory stream and also append newly generated memories back into it.
Action Execution Module.
Appendix B A/B Testing Scenario
@@
We designed a simulated A/B testing scenario focused on a novel feature design of the left-side filtering panel of Amazon’s shopping interface, a common area for UI experimentation in e-commerce. Our team has the privilege of altering the left filter panel interface (Fig. 7). In comparison to the existing design as the control condition, where all the filter options are shown, the new design (treatment condition) uses a similarity-based ranking algorithm to reduce the filter options that have a lower-than-80% similarity score to the user’s search query.
@@
Using the LLM Agent Generation Module, we created 100,000 agent personas [^3] We will open source these agent personas and then randomly sampled 1,000 of them to simulate individual shopping sessions, with 500 assigned to the treatment condition and 500 to the control condition. The personas’ demographic attributes were sampled to match real online shopper distributions, including age (18–24: 15.8%, 25–34: 24.6%, 35–44: 22.1%, 45–54: 19.6%, 55–64: 17.9%), gender (49.4% male, 50.6% female), and household income (0–30,000: 32.1%, 30,001–94,000: 32.4%, 94,001–1,000,000: 35.5%). In parallel, we also conducted the real human A/B testing with 2 million users randomly assigned into the control condition and the treatment condition. Each agent was initialized with a persona profile and a shopping task (e.g., “find a budget smart speaker under \$40 with strong customer reviews”). The allowed action space for the LLM agent includes typical e-commerce flows: search, filter use, product examination, and cart actions. Persona generation followed the methodology in Chen et al. (2024) with LLM agent’s demographics (age, income, occupation), preferences, and shopping goals.
@@
To learn the environment cost of LLM Agents (as suggested by (Inie et al., 2025)), we calculated the number of tokens and CO2 emissions of our experiment. For 1,000 agent simulations using Claude 3.5 Sonnet, estimated 875 million tokens, resulting in a cost of approximately \$2,925 and CO2 emissions of 35 kilograms (using 0.00004 g CO2/token for average data center grid mix). In comparison, recruiting 1,000 human participants into a typical UX study session in the AI Industry would cost \$100 per person (\$100,000 in total).
Figure 5. Number of LLM agents who completed a purchase under control and treatment conditions.
A bar chart comparing the number of LLM agents who made a purchase. The control group with all filter options has 403 purchases, while the treatment group with reduced filter options has 414 purchases.